Calculate the consumption of goods and services in your family in the following categories:

a. Groceries like food items Consumption Prices Amount(Rs.)

(A) Meat items

1. Chicken

2. Beef

3. Mutton

4. Any other

b. Pulses

c. Spices

d. Other items

e. Utilities

(A) Electricity

(B) Gas

(C) Water

(D) Internet

(E) Drainage and Conservancy

(F) Any Other

f. Conveyance

(A) Father’s office

(B) Mother’s shopping

(C) Siblings schooling

(D) Yours University

(E) Any other (if any)

g. Tuition Fees (if any)

(A) School…(Siblings)

(B) College..(Siblings)

(C) University (Yours + Siblings)

(D) Tuition (Yours + Siblings)

(E) Any Other

h. Medical Charges (if any)

(A) Grand Parents (if any)

(B) Parents (if any)

(C) Yours and Siblings (if any)

(D) Any other


1
Expert's answer
2021-04-21T10:57:41-0400

Consumption in economics is the total value of goods and services consumed by an consumer. Symbolically it is shown as

C = a + b Yd

a = autonomous Consumption

b = Slope

Yd= Disposable Income

 

(a.meat items

1. Chicken=500

2. Beef=1000

3. Mutton=200

Consumption = Chicken + Beef + Mutton

"=50 + 30 + 20"

"=100" Rs

b. Pulses=50

Consumption=pulses

"=50" Rs

c. Spices=50

Consumption=spices

"=50" Rs

e. Utilities

(A) Electricity=25

(B) Gas=30

(C) Water=15

(D) Internet=10

(E) Drainage and Conservancy=10

Consumption=electricity+gas+water+internet+drainage

"=25+30+15+10+10=90" Rs

f. Conveyance

(A) Father’s office=30

(B) Mother’s shopping=100

(C) Siblings schooling=100

(D) Yours University=400

Consumption=father's office + mother's shopping+sibling schooling+university

"=30+100+100+400=630" Rs

g. Tuition Fees (if any)

(A) School…(Siblings)=20

(B) College..(Siblings)=500

(C) University (Yours + Siblings)=200

(D) Tuition (Yours + Siblings)=300

Consumption=school+college+university+tuition

"=20+500+200+300=1020" Rs

h. Medical Charges (if any)

(A) Grand Parents (if any)=0

(B) Parents (if any)=100

(C) Yours and Siblings (if any)=200

Consumption=grandparents+parents+siblings

"=0+100+200=300" Rs

Total family consumption=meat+pulses+spices+utilities+conveyance+tuition fees+medical charges"=100+50+50+90+630+1020+300"

"=2240" Rs